OpenStudy (anonymous):

Knowing that A^-1=b+c and C=(a+b)^-1 Demonstrate AB=BC

OpenStudy (kainui):

Gonna need a little more information than that. Are A and C similar matrices?

ganeshie8 (ganeshie8):

http://www.allmathwords.org/images/2m3dot.gif

OpenStudy (gorv):

\[a^{-1}=\frac{ 1 }{ a }\]

OpenStudy (gorv):

\[\frac{ 1 }{ a}=b+c\]

OpenStudy (gorv):

\[c=\frac{ 1 }{ a+b}\]

OpenStudy (gorv):

\[a+b=\frac{ 1 }{ c}\]

OpenStudy (gorv):

ab+ac=1 ac+bc=1 ab+ac=ac+bc ab=bc

OpenStudy (gorv):

thatzzz it

OpenStudy (gorv):

@Pepita

OpenStudy (gorv):

the info u give is enough to solve the Q